OpenStudy (anonymous):

–3x + 2y = 8 Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (isaiah.feynman):

First re write the equation in the form y=mx+b. We know the line will be straight, so the line will be stretched by a factor of m, then shifted by c units upwards. :)

OpenStudy (anonymous):

i got y=3/2x+8 but what are the plots?

OpenStudy (isaiah.feynman):

Yes, I was working hard on that few minutes ago. I got (0,4) and (-8/3,0) I have verified those plots by using them to find the slope, and I got 3/2 as slope!!! So those are your plots!!! :D

OpenStudy (anonymous):

thanks and how would i write these plots in a complete sentence explaining how to graph them?

OpenStudy (isaiah.feynman):

Umm... you can say you simply plotted the points on the line. If you want to know how I got the points, I made x=0 then I got y=4 thus (0,4). Again I made y=0 then I got x=-8/3 thus (-8/3,0).